Please help me solve this equation!
The product of two consecutive odd integers is 323. Find the numbers.

Answer by oscargut(2103)

you have to resolve x(x+2)=323 and select the odd solution so:
x^2+2x-323=0 x=(-2+36)/2 or x=(-2-36)/2 so solutions are 17 and -19 so x=17 and x+2=19 so the numbers are 17 and 19